$SPAD/src/input c02aff.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
a:Matrix SF:=
 [[5.0 ,30.0 ,-0.2 ,50.0 ,-2.0 ,10.0 ],_
  [6.0 ,20.0 ,-6.0 ,100000.0 ,-40.0 ,1.0 ]]
n:=ncols(a)-1
scale:=true
result:=c02aff(a,n,scale,-1)

———-

References^{[1] nothing}

4